OpenStudy (anonymous):

Find the equation for the normal line to f at the point (-4, f(-4)). f(x)=-2x+3x^2

OpenStudy (anonymous):

find the derivative. plug in -4 to get the slope normal means line is perpendicular so slope is the negative reciprocal find the equation for the line with that slope through the point given

OpenStudy (anonymous):

want to do it?

OpenStudy (amistre64):

y = \(-\frac{1}{f'(x)}(x+4)+f(x)\)

OpenStudy (amistre64):

i spose f'(-4) and f(-4) are better markers in it :)

OpenStudy (anonymous):

guru ! should be \[+f(-4)\] why am i always cleaning up after...

OpenStudy (amistre64):

i caught it :)

OpenStudy (anonymous):

i know i just wanted to crow. back to the dungeon...

OpenStudy (anonymous):

haha! what is the question is telling you to take the derivative? is the line a tangent?

OpenStudy (anonymous):

you got this or you need help?

OpenStudy (anonymous):

no line is not tangent. "normal" means perpendicular

OpenStudy (anonymous):

lets do it.

OpenStudy (anonymous):

\[f(x)=3x^2-2x\] \[f(-4)=56\] so the point is (-4,56)

OpenStudy (anonymous):

ok, gotcha

OpenStudy (anonymous):

\[f'(x)=6x-2\] \[f'(-4)=-26\] so slope of tangent line is -36

OpenStudy (anonymous):

slope of "normal" line is \[\frac{1}{26}\]

OpenStudy (anonymous):

you have the slope, you have the point, so find the equation of the line via \[y-56=\frac{1}{26}(x+4)\]

OpenStudy (anonymous):

or you can write \[y-56=\frac{1}{26}x+\frac{2}{3}\]

OpenStudy (anonymous):

one sec, Im going through the steps

OpenStudy (anonymous):

ok i will wait. it is now a problem of finding the equation of a line given a slope and a point

OpenStudy (amistre64):

\[y = -\frac{1}{f'(-4)}(x+4)+f(-4)\]

OpenStudy (anonymous):

ok, so the derivative is used to calc the slope?

OpenStudy (anonymous):

yes. the derivative itself is not a slope, but it is the formula for one. so for example you know the derivative of \[f(x)=x^2\] is\[f'(x)=2x\] this means the slope of the line tangent to \[y=x^2\] at (3,9) is 6 etc

OpenStudy (anonymous):

ok, gotcha

OpenStudy (anonymous):

but of course in this case finding the slope was one step on the way to finding the slope of the perpendicular line

OpenStudy (anonymous):

we got a slope of -26 so the perpendicular line has slope \[\frac{1}{26}\]

OpenStudy (anonymous):

understand- thanks, I just did another one

OpenStudy (anonymous):

good!

OpenStudy (anonymous):

where does the (x+4) come from> In my country we always finished with y=mx+c

OpenStudy (anonymous):

thank!

OpenStudy (anonymous):

nothing comes up

OpenStudy (anonymous):

trying to find the example we did earlier on today where we had x-4/absolute(x-4) and we had to clac left and and right hand side where x=-4

OpenStudy (amistre64):

there are a lot of different 'forms' that an equation of a line can take; y = mx +c is just one of them

OpenStudy (amistre64):

y = mx -mx0 + y0 is another where your 'c' values = (-mx0+y0)